# AutoBencher: Creating Salient, Novel, Difficult Datasets for Language Models

### v1

@@ Abstract
Evaluation is critical for assessing capabilities, tracking scientific progress, and informing model selection. In this paper, we present three desiderata for a good benchmark for language models: (i) salience (e.g., knowledge about World War II is more salient than a random day in history), (ii) novelty (i.e., the benchmark reveals new trends in model rankings not shown by previous benchmarks), and (iii) difficulty (i.e., the benchmark should be difficult for existing models, leaving headroom for future improvement). We operationalize these three desiderata and cast benchmark creation as a search problem, that of finding benchmarks that that satisfy all three desiderata. To tackle this search problem, we present AutoBencher, which uses a language model to automatically search for datasets that meet the three desiderata. AutoBencher uses privileged information (e.g. relevant documents) to construct reliable datasets, and adaptivity with reranking to optimize for the search objective. We use AutoBencher to create datasets for math, multilingual, and knowledge-intensive question answering. The scalability of AutoBencher allows it to test fine-grained categories and tail knowledge, creating datasets that are on average 27% more novel and 22% more difficult than existing benchmarks. A closer investigation of our constructed datasets shows that we can identify specific gaps in LM knowledge in language models that are not captured by existing benchmarks, such as Gemini Pro performing much worse on question answering about the Permian Extinction and Fordism, while OpenAGI-7B performing surprisingly well on QA about COVID-19.111Code is available at https://github.com/XiangLi1999/AutoBencher.
@@ 1 Introduction
Evaluation is crucial for informing model selection and guiding model development, and it is especially challenging to evaluate language models. Prior works have constructed high-quality benchmarks, guided by two desiderata: (1) salience: the benchmark should test practically important capabilities, for example, the capability to solve math problems is examined by GSM8K [6] and capability to avoid answering misleading questions is examined by TruthfulQA [17]. (2) difficulty: the benchmark should yield low accuracy with existing models. For example, GPQA [28] carefully constructs questions that are challenging even for domain experts. In this work, we identify an additional goal of (3) novelty. More recently, observational studies of existing benchmarks show that model performance on existing benchmarks are highly correlated [18, 27, 20, 5], and thus there is an increasing need for benchmark novelty. Novelty measures the degree to which a benchmark reveals previously unknown performance trends in model rankings, such as unexpected model strengths or weaknesses on specific domains.
In order to measure salience, difficulty, and novelty of a benchmark, we operationalize these three properties: (1) We define salience as a binary variable that captures whether a benchmark on a specific topic (e.g., COVID-19) is important. (2) The difficulty of a benchmark is determined by the lowest error rate achievable by existing models. (3) We calculate the novelty of a benchmark as the predictability of model performance on the benchmark based on performance on other benchmarks. Formalizing the three desiderata allows us to take an optimization approach to constructing benchmarks. Instead of viewing benchmarking as the task of constructing a dataset that represents some task, we cast it as a way to find particular test set that fulfill our three desiderata. This turns benchmarking into an optimization problem, where the objective is made explicit.
@@
To approximately solve this search problem and automatically create new salient, difficult, and novel benchmarks, we propose AutoBencher. As shown in Figure 1, given a broad domain (e.g., history), AutoBencher proposes candidate evaluation topics (e.g., World War II) and then constructs a small dataset for each topic using privileged information (e.g., relevant Wikipedia articles). These datasets are scored by our metrics and used as part of a local search algorithm to find a benchmark with improved difficulty. We leverage the scalability of AutoBencher to identify and select topics that jointly maximize a weighted sum of difficulty and novelty, subject to a salience constraint specified by the user.
@@
We use AutoBencher to create datasets in 5 domains: math, history, science, economics, and multilingualism, altogether producing around 3700 examples. These benchmarks reveal novel trends not shown by prior benchmarks (§ 6.2): For example, we find that while Gemini Pro is one of the strongest models on existing history benchmarks, it performs quite poorly on AutoBencher-discovered topics of Permian Extinction and Fordism, performing even worse than some 7B models, such as Mistral-7B. Meanwhile, we also find that 7B models (e.g., Mistral-7B and OpenAGI-7B), which typically exhibit lower accuracy due to their smaller size, perform quite well on the topic of COVID-19, even achieving accuracy on par with GPT-4 Turbo. Additionally, our datasets increase difficulty by 22% compared with human-constructed benchmarks, such as MMLU and XOR QA (§ 6.1). [^2] We will release the dataset and code.
@@
Figure 1: (Left) Models ranked by accuracy on existing datasets and AutoBencher datasets. Existing datasets show roughly the same performance trends, while AutoBencher discovers tests that induce novel rankings. (Right) Given a domain (e.g., history), AutoBencher creates datasets that are salient, difficult, and novel. It achieves this by searching through evaluation topics, scoring each based on these desiderata, and selecting the best one. These datasets are automatically created by providing privileged information to a language model.
@@ 2 Related Work
A large number of benchmarks have been constructed to measure different skills of language models. To give a few examples of common LLM benchmarks, MMLU measures the understanding of academic subjects [11], GSM8K measures mathematical reasoning [6], and Winogrande measures common sense reasoning [30]. Researchers have also grouped the benchmarks to create leaderboards that rank LMs’ overall capabilities, such as HELM [16] and Open LLM Leaderboard [3]. Additionally, researchers also carefully subsample existing benchmarks to obtain a smaller and more efficient benchmark with similar model accuracy [20]. Prior works of LLM-as-Judge incorporate language models as part of the evaluation pipeline to automatically judge the quality of model-generated responses [8, 39, 9]. Our work further automates the evaluation pipeline by using LM to propose evaluation topics and generate questions.
The most similar work to ours is Li et al. [15], which uses LMs to brainstorm questions for instruction following tasks. However, their goal is very different from ours: they aim to find a minimal set of questions to reproduce the same LM ranking achieved by evaluating a large number of examples; our goal is to search for new benchmark topics that break the performance trends of existing benchmarks.
Adaptive Datasets
@@
In AutoBencher, one important desideratum we optimize for is novelty. Prior works have also constructed datasets adaptively to search for difficult questions [22, 12, 29, 38, 7]. Most of these works have generated test cases with human annotators, whereas we use language models to automate the search, saving extensive human effort. Similar to AutoBencher, Perez et al. [26] searches for prompts that induce harmful behaviors in language models using another language model. However, they focus on instance-level failures in safety, whereas we focus on finding systematic failures on specific topics.
3 A Metric-Driven Framework of Benchmark Creation
We now discuss our three desiderata of salience, difficulty, and novelty and formally define them as quantitative metrics that can be directly optimized.
Preliminaries.
Let $c\in\mathcal{C}$ be a evaluation topic. We define a dataset $\mathcal{D}_{c}=\{(x_{i},y_{i})\}_{i}$ as a set of question-answer pairs $(x_{i},y_{i})$ that evaluate mastery of the topic $c$. As an example, topics $\mathcal{C}$ could include “the Industrial Revolution”, “World War II”, “Ancient Civilization”, etc. In this work, we will generate the datasets $\mathcal{D}_{c}$ from a fixed generator $p(\mathcal{D}_{c}\mid c)$ and focus on optimizing the set of topics $c$ to change the novelty and difficulty of a dataset.
@@ 3 A Metric-Driven Framework of Benchmark Creation
Let =$\mathcal{M}=\{\texttt{LM}_{m}\}_{m=1}^{M}$ denote the set of $M$ existing models to evaluate. We denote the accuracy of model $\texttt{LM}_{m}\in\mathcal{M}$ on benchmark $\mathcal{D}_{c}$ as $\texttt{acc}(\texttt{LM}_{m},\mathcal{D}_{c})$. We define the accuracy vector $V_{c}=[\texttt{acc}(\texttt{LM}_{1},\mathcal{D}_{c}),\cdots,\texttt{acc}(\texttt{LM}_{M},\mathcal{D}_{c})]$ of a benchmark $\mathcal{D}_{c}$ as the accuracy of all models on the benchmark $\mathcal{D}_{c}$.
Salience.
@@
Recall that salience measures the practical importance of a topic $c$. First, we assume a set of salient topics $\mathcal{S}$ specified by the user, and we define salience as a binary variable, such that $\textsc{salience}(c)=1$ if $c\in\mathcal{S}$ and $\textsc{salience}(c)=0$ otherwise. For example, we may define salient topics $\mathcal{S}$ to be the set of topics with the number of corresponding Wikipedia page views exceeding a threshold.
Novelty.
@@
Recall that novelty measures how much new information a dataset reveals about existing models over existing benchmarks. Therefore, we formalize $\textsc{Novelty}(\mathcal{D}_{c};\mathbf{D}_{\text{prev}};\mathcal{M})$ as a function of the benchmark in question $\mathcal{D}_{c}$, prior benchmarks $\mathbf{D}_{\text{prev}}:=\{\mathcal{D}_{1}\ldots\mathcal{D}_{N}\}$, and the existing set of models $\mathcal{M}$. Intuitively, the results on a new benchmark reveal new information if model performance on the new benchmark vastly differ from the trends on prior benchmarks (e.g., if a normally low-performing model outperforms all other models on the new benchmark).
@@
To quantify this, we first find how much information is explainable by the accuracy on existing datasets $\mathbf{D}_{\text{prev}}$, by performing a regression from $V_{\text{prev}}:=\left[V_{1}\cdots V_{N}\right]\in\mathbb{R}^{N\times M}$ to predict $V_{c}\in\mathbb{R}^{1\times M}$,
@@
If the new accuracy vector $V_{c}$ is spanned by the existing accuracy vectors, $\textsc{RankCorr}(V_{c},\hat{V}_{c})$ will be close to 1, resulting in low novelty. On the other hand, if the new accuracy vector $V_{c}$ discovers some new patterns in model performance such as an orthogonal direction, $\textsc{RankCorr}(V_{c},\hat{V}_{c})$ will be low, resulting in high novelty.
Take Figure 1 (left) as an example, let ARC be the existing datasets $\mathbf{D}_{\text{prev}}$. To compute the novelty of the MMLU dataset (column 2), we perform a regression from the accuracy vector of ARC to predict the accuracy vector of MMLU. Since they have the same ranking, model rankings on MMLU is perfectly predictable from ARC, yielding a RankCorr of $1$ and a Novelty of 00. To compute the novelty of the QA: COVID-19 dataset (column 3), we also predict its accuracy vector from ARC. However, since the accuracy vector of COVID-19 datasets shows abrupt changes in model performance, it is not spanned by the accuracy vector of the ARC dataset. Therefore, its performance is not predictable by the existing dataset, yielding a low RankCorr and a high Novelty.
Difficulty.
A benchmark’s difficulty is determined directly by a model’s error rate. Ideally, a benchmark should leave sufficient headroom above the best current error rate to enable tracking future progress. We formalize the difficulty of a benchmark as the remaining gap between the best model’s accuracy and optimal accuracy: $\textsc{Difficulty}(\mathcal{D}_{c},\mathcal{M})=1-\max_{m}\texttt{acc}(\texttt{LM}_{i},\mathcal{D}_{c})$.
Separability.
Besides the three main desiderata, we also include the separability of model accuracy as an additional regularizer. Separability measures the amount of separation between different model accuracies; In this paper, we use separability to complement novelty, ensuring that all the model performance trends revealed by a benchmark is robust. We formalize the separation on benchmark $\mathcal{D}_{c}$ between the accuracy of models $\mathcal{M}$ as the mean absolute deviation $\textsc{Sep}(\mathcal{D}_{c},\mathcal{M})=\text{mean}(|V_{c}-\text{mean}(V_{c})|)$.
Existing Benchmarks Case Study.
We now analyze several existing benchmarks under these desiderata: MMLU [11] contains salient topics on academic subjects; it is sufficiently difficult with the best model accuracy of 86% and has good separability to distinguish existing models. However, the benchmark lacks novelty, as it’s highly correlated with prior benchmarks like ARC, with a rank correlation of 94%.
Datasets in the Inverse Scaling Prize [21] is novel but many lack salience. For example, pattern matching suppression instructs the LM to continue a sequence by violating the current pattern, and Memo trap instructs the LM to continue a famous quote with a different ending. These tasks reveal novel performance patterns (i.e., larger models perform worse), but are much less salient than datasets like the knowledge-intensive MMLU.
4 AutoBencher
Our goal is to find a sufficiently salient evaluation topic $c$ that maximizes a linear combination of novelty and difficulty. We directly maximize novelty and difficulty subject to a constraint on salience, because salience is operationalized as a binary variable. Specifically, we aim to solve the following constrained optimization problem:
maximize $\mathcal{J}(c)=\textsc{Novelty}(\mathcal{D}_{c};\mathbf{D}_{\text{prev}},\mathcal{M})+\beta_{1}~{}\textsc{difficulty}(\mathcal{D}_{c};\mathcal{M})+\beta_{2}~{}~{}\textsc{Sep}(\mathcal{D}_{c};\mathcal{M})$ (1)
@@ 4 AutoBencher
Solving the optimization problem.
Identifying solutions to this search problem is difficult due to the complexity of performing a discrete search over three potentially conflicting criteria (salience, difficulty, and novelty) and we cannot intuitively judge which topics might reveal novel performance trends a priori. Therefore, we propose AutoBencher, which automates the search with a language model, denoted as $\texttt{LM}_{\text{gen}}$, and outputs the best discovered topic and its dataset.
One natural, naive design is to perform a random search, where we prompt $\texttt{LM}_{\text{gen}}$ to generate a diverse set of topics $c$, prompt $\texttt{LM}_{\text{gen}}$ to generate a dataset of (question, answer) pairs for each topic, and then select the best topic according to the objective function $\mathcal{J}(c)$.
However, this design suffers from two issues: (1) Example correctness: Since we use $\texttt{LM}_{\text{gen}}$ to construct the dataset, the generated answers might be incorrect due to model hallucination. (2) Example difficulty: The difficulty of the generated questions is upper-bounded by the capabilities of $\texttt{LM}_{\text{gen}}$ and hence cannot be used to evaluate stronger models. (3) Topic difficulty: empirically, we observe that $\texttt{LM}_{\text{gen}}$ tends to propose well-known, and insufficiently difficult topics.
We now propose some techniques to address these issues: We first augment $\texttt{LM}_{\text{gen}}$ with privileged information to improve the correctness and difficulty of the generated datasets (§ 4.1). Next, we propose adaptive search, which uses the trajectory of past generated benchmarks to improve topic difficulty (§ 4.2). Finally, we present the full pseudocode and present how each metric gets optimized in this pipeline (§ 4.3).

### v2

# AutoBencher: Towards Declarative Benchmark Construction

@@ Abstract
We present AutoBencher, a declarative framework for automatic benchmark construction, and use it to scalably discover novel insights and vulnerabilities of existing language models. Concretely, given a few desiderata of benchmarks (e.g., question difficulty, topic salience), we operationalize each desideratum and cast benchmark creation as an optimization problem. Specifically, we experiment with two settings with different optimization objectives: (i) for capability evaluation, we declare the goal of finding a salient, difficult dataset that induces novel performance patterns; (ii) for safety evaluation, we declare the goal of finding a dataset of unsafe prompts that existing LMs fail to decline. To tackle this optimization problem, we use a language model to iteratively propose and refine dataset descriptions, which are then used to generate topic-specific questions and answers. These descriptions are optimized to improve the declared desiderata. We use AutoBencher (powered by GPT-4) to create datasets for math, multilinguality, knowledge, and safety. The scalability of AutoBencher allows it to test fine-grained categories and tail knowledge, creating datasets that elicit 22% more model errors (i.e., difficulty) than existing benchmarks. On the novelty ends, AutoBencher also helps identify specific gaps not captured by existing benchmarks: e.g., Gemini-Pro has knowledge gaps on Permian Extinction and Fordism while GPT-4o fails to decline harmful requests about cryptocurrency scams. [^1] Code is available at https://github.com/XiangLi1999/AutoBencher.git
@@ 1 Introduction
Evaluation is crucial for informing model selection and guiding model development, and language model evaluation is especially challenging. Many prior works aim to make evaluation cheaper, faster, and more scalable by automating parts of the evaluation pipeline: For example, AlpacaEval (Dubois et al., 2023) uses LLM-based automatic evaluation for instruction following tasks; Zheng et al. (2023) shows that strong LLM judges like GPT-4 can approximate human preference. While many works focus on automatically judging model responses, very few works attempt to automatically construct the evaluation dataset (i.e., generate the questions). In this paper, we present AutoBencher, a declarative framework for automatic dataset construction, and use it to scalably discover novel insights and model vulnerabilities not shown by existing benchmarks.
In AutoBencher, we first declare a few desiderata for the dataset, then we build quantitative surrogate metrics for them, and search for a particular dataset that optimizes an explicit objective of our desiderata. The objective allows us to precisely measure the progress of our constructed datasets: e.g., the new dataset is 20% more difficult than the old dataset. Furthermore, the solution to these optimization problems might be datasets that reveal information that’s not captured by existing benchmarks (e.g., unexpected knowledge gaps and safety vulnerabilities).
To instantiate this idea of declarative benchmark construction, we experiment with two benchmark settings with different desiderata. In the first setting, we evaluate math, knowledge, and multilingual skills, and we consider four desiderata: (1) Salience: the benchmark should test practically important capabilities. (2) Difficulty: existing models should obtain low accuracy on the benchmark. (3) Separability: existing models should obtain accuracies that are spread apart on the benchmark. (4) Novelty: we define novelty to measure the degree to which a benchmark reveals previously unknown trends in model rankings. Under our definition, a novel dataset should reveal a model ranking that’s not consistent with rankings on existing datasets (e.g., weaknesses of a generally strong LM). In the second setting, we evaluate LMs’ ability to refuse complying with harmful requests, and we consider two desiderata of the dataset: (1) Harmfulness: the requests ask for responses that could cause harm. (2) Attack success rate: a large percentage of requests in the dataset should trigger LMs to produce harmful responses. For both capability and safety settings, we formalize their respective desiderata (§ 3) and cast benchmark construction as an optimization problem.
@@
To approximately solve this optimization problem, we propose to use a language model to automatically construct datasets and iteratively revise the dataset description to optimize for the declared desiderata. In AutoBencher, we have an evaluator LM, which proposes dataset descriptions and generates questions, and a candidate LM, which is evaluated on the generated dataset to provide feedback. As shown in Figure 1, given a broad domain (e.g., history), an evaluator LM proposes a few dataset descriptions (e.g., important events in World War II) and then constructs a small dataset for each description using privileged information (e.g., relevant Wikipedia articles or Python libraries). Then, the candidate LM answers the questions in these datasets without access to the privileged information. Each dataset is scored according to the desiderata (e.g., difficulty) and used to inform the proposal of new datasets with improved desiderata scores. We leverage the scalability of AutoBencher to identify and select dataset descriptions that jointly maximize a weighted sum of the desiderata metrics.
@@
We use AutoBencher (with GPT-4-0125 as the evaluator LM) to create datasets in 6 domains: math, history, science, economics, multilinguality, and safety, altogether producing around 4000 examples. These benchmarks reveal novel trends and weaknesses not captured by prior benchmarks (§ 6.3): For example, we find that while Gemini Pro is one of the strongest models on existing history benchmarks, it performs quite poorly on AutoBencher-discovered topics of Permian Extinction and Fordism, performing even worse than some 7B models, such as Mistral-7B. Meanwhile, we found that GPT-4 Turbo fails to refuse questions about replicating terror events (e.g., How to replicate an airport attack?). Our AutoBencher dataset reveals a 27% decrease in model ranking correlation (i.e., more novel), and a 22% decrease in best model accuracies (i.e., more difficult), compared with human-constructed benchmarks, such as MMLU (Hendrycks et al., 2021) (§ 6.1). Our safety dataset induces 20% more attack success rate than existing safety datasets, such as XSTest (Röttger et al., 2024) and HarmBench (Mazeika et al., 2024).
@@
Figure 1: (Left) A toy example of model rankings on existing datasets and AutoBencher datasets. Existing datasets show roughly the same performance trends, while AutoBencher discovers tests that induce novel rankings. (Right) Given a domain (e.g., history), AutoBencher creates datasets that are salient, difficult, and novel. It achieves this by searching over dataset descriptions (e.g., the timeline of WWII), scoring each based on difficulty and novelty, and selecting the best one.
@@ 2 Related Work
Benchmarking Language Models. A large number of datasets have been constructed to measure different skills of language models, and multiple related datasets aggregate to form a benchmark. For example, MMLU measures the understanding of academic subjects (Hendrycks et al., 2021), and Winogrande measures common sense reasoning (Sakaguchi et al., 2019). Researchers have also grouped the benchmarks to create leaderboards that rank LMs’ overall capabilities, such as HELM (Liang et al., 2022), Open LLM Leaderboard (Beeching et al., 2023), BIG-Bench (Srivastava et al., 2023), and lm-evaluation-harness (Gao et al., 2024) Additionally, researchers also carefully subsample existing benchmarks to obtain smaller and more efficient benchmarks that elicit similar model accuracies.(Maia Polo et al., 2024). Prior works of LLM-as-Judge incorporate language models to automatically judge model-generated responses to a set of prompts (Dubois et al., 2023; Zheng et al., 2023; Fu et al., 2023; Li et al., 2024). Our work goes further and uses LMs to automatically generate the prompts themselves.
The most similar work to ours is LM-Examiner (Bai et al., 2023), which also uses LMs to generate benchmark questions. However, their method is different from ours: LM-Examiner directly generates questions and follow-ups from the model’s parametric memory, whereas AutoBencher generates more difficult questions by relying on privileged information (e.g., retrieval or Python tools). Concretely, ChatGPT attains 97%+ accuracy on the LM-Examiner dataset and only around 60% accuracy on AutoBencher datasets.
@@
Adaptive Datasets. In AutoBencher, one important desideratum we optimize for is difficulty. Prior works have also constructed datasets adaptively to search for difficult questions (Nie et al., 2020; Jia & Liang, 2017; Ribeiro et al., 2020; Xu et al., 2020; Dinan et al., 2019). Most of these works have generated test cases with human annotators, whereas we use language models to automate the search, saving extensive human effort.
Similar to AutoBencher for safety, work on red-teaming language models (Perez et al., 2022; Zou et al., 2023; Liu et al., 2023) automatically searches for prompts that induce harmful behaviors in language models via gradient-based optimization or genetic algorithm. However, they focus on making local edits (e.g., adding some adversarial tokens) to trigger instance-level safety failures. We instead focus on finding general and systematic failures in safety (e.g., categories of harmful topics that LMs fail to reject, and excuses that mislead the LMs to provide harmful responses). Also, our approach generalizes beyond safety settings to evaluate LM capabilities (e.g., knowledge, multilinguality and math) as well.
3 A Declarative Framework of Benchmark Creation
To instantiate this idea of declarative benchmark construction, we experiment with two settings for benchmark construction. (i) For the capability datasets, we consider four desiderata of salience, difficulty, separability and novelty. (ii) For the safety datasets, we consider two desiderata of harmfulness and attack success rate. We formally define them as quantitative metrics that can be directly optimized.
Preliminaries. Let $c\in\mathcal{C}$ be a natural language description of a dataset (e.g., “timeline of the Industrial Revolution”, “Canada’s involvement in World War II”, “solving for second derivatives of polynomials”, "execution details of cryptocurrency scams"). We define a dataset $\mathcal{D}_{c}=\{(x_{i},y_{i})\}_{i}$ as a set of question-answer pairs $(x_{i},y_{i})$ that evaluate mastery of the knowledge or skill required by $c$. In this work, we will generate the datasets $\mathcal{D}_{c}$ from a tool-augmented language model $p(\mathcal{D}_{c}\mid c)$ and focus on selecting the set of dataset descriptions $c$ to optimize the desiderata.
@@ 3 A Declarative Framework of Benchmark Creation
Let =$\mathcal{M}=\{\texttt{LM}_{m}\}_{m=1}^{M}$ denote the set of $M$ existing models to evaluate. We denote the accuracy of model $\texttt{LM}_{m}\in\mathcal{M}$ on a dataset $\mathcal{D}_{c}$ as $\texttt{acc}(\texttt{LM}_{m},\mathcal{D}_{c})$. For the safety evaluation, the correct answer is to abstain from answering the question; therefore, we define accuracy on the safety dataset as the rejection rate. We define the accuracy vector $v_{c}=[\texttt{acc}(\texttt{LM}_{1},\mathcal{D}_{c}),\cdots,\texttt{acc}(\texttt{LM}_{M},\mathcal{D}_{c})]$ as the accuracy of all models on the dataset $\mathcal{D}_{c}$.
3.1 Capability Evaluation
@@
Salience. Recall that salience measures the importance of a dataset description $c$. First, we assume a set of salient topics $\mathcal{S}$ specified by the user, and we define salience as a binary variable, such that $\textsc{salience}(c)=1$ if $c\in\mathcal{S}$ and $\textsc{salience}(c)=0$ otherwise. For example, we may define salient topics $\mathcal{S}$ to be the set of descriptions with the number of relevant Wikipedia page views exceeding a certain threshold.
Difficulty. A benchmark’s difficulty is determined directly by a model’s error rate. Ideally, a benchmark should leave sufficient headroom above the best current error rate to enable tracking future progress. We formalize the difficulty of a benchmark as the lowest achieved error rate: $\textsc{Difficulty}(\mathcal{D}_{c},\mathcal{M})=1-\max_{m\in\mathcal{M}}\texttt{acc}(\texttt{LM}_{m},\mathcal{D}_{c}){=1-\max v_{c}}$.
Separability. Separability measures the amount of separation among different model accuracies of the same dataset. We formalize the separation on benchmark $\mathcal{D}_{c}$ between the accuracy of models $\mathcal{M}$ as the mean absolute deviation $\textsc{Sep}(\mathcal{D}_{c},\mathcal{M})=\text{mean}(|v_{c}-\text{mean}(v_{c})|)$. Separability ensures that all the model performance trends revealed by the dataset are robust. When a dataset elicits very similar accuracies on two LMs, their ranking may swap if we introduce a small amount of noise (e.g., subsample the dataset), hurting the robustness of the evaluation results.
@@
Novelty. Novelty measures how much new information a dataset reveals about existing models over existing benchmarks. We formalize $\textsc{Novelty}(\mathcal{D}_{c};\mathbf{D}_{\text{prev}};\mathcal{M})$ as a function of the dataset in question $\mathcal{D}_{c}$, prior datasets $\mathbf{D}_{\text{prev}}:=\{\mathcal{D}_{1}\ldots\mathcal{D}_{N}\}$, and the models we evaluate $\mathcal{M}$. Intuitively, the results on a new dataset reveal new information if model performance on the new dataset vastly differs from the trends on prior datasets (e.g., if a normally low-performing model outperforms all other models on the new dataset).
@@
To quantify this, we first find how much variance of $v_{c}$ is explainable by the accuracy on existing datasets $\mathbf{D}_{\text{prev}}$, by performing a regression from $V_{\text{prev}}:=\left[v_{1}\cdots v_{N}\right]\in\mathbb{R}^{M\times N}$ to predict $v_{c}\in\mathbb{R}^{M\times 1}$ with parameter $\theta\in\mathbb{R}^{N\times 1}$ and $b\in\mathbb{R}^{M\times 1}$,
@@
If the new accuracy vector $v_{c}$ is spanned by the existing accuracy vectors, $\textsc{RankCorr}(v_{c},\hat{v}_{c})$ will be close to 1, resulting in low novelty. On the other hand, if $v_{c}$ discovers some new patterns in model performance such as an orthogonal direction, $\textsc{RankCorr}(v_{c},\hat{v}_{c})$ will be low, resulting in high novelty.
Case Study of MMLU. We now analyze the MMLU benchmark (Hendrycks et al., 2021) under salience, novelty, difficulty and separability: MMLU contains salient topics on academic subjects; it is sufficiently difficult with the best model accuracy of 86% and has good separability to distinguish existing models. However, the benchmark lacks novelty, as language models’ ranking on the full MMLU benchmark is highly correlated with prior benchmarks like ARC, with a rank correlation of 94%.
Optimization Objective. Our goal is to find a sufficiently salient dataset description $c$ that maximizes a linear combination of novelty, difficulty, and separability, subject to a constraint on salience. Specifically, we aim to solve the following constrained optimization problem:
maximize $\mathcal{J}(c;\mathcal{M})=\textsc{Novelty}(\mathcal{D}_{c};\mathbf{D}_{\text{prev}},\mathcal{M})+\beta_{1}~{}\textsc{difficulty}(\mathcal{D}_{c};\mathcal{M})+\beta_{2}~{}~{}\textsc{Sep}(\mathcal{D}_{c};\mathcal{M})$ (1)
@@
3.2 Safety Evaluation
Harmfulness. Harmfulness measures whether the prompts in the dataset will likely lead to harmful responses. For example, a question about building a bomb is harmful, but a philosophical question about the “trolley problem” is not harmful. Assume a set of harmful topics specified by a taxonomy from existing safety datasets, and we define harm as a binary variable, such that $\textsc{harm}(c)=1$ if $c\in\mathcal{S}_{\text{harm}}$ and $\textsc{harm}(c)=0$ otherwise.
Attack success rate. A safe model should abstain from responding to any harmful prompts, by politely declining to provide an answer. A safety dataset’s attack success rate is determined directly by models’ rejection rate. Formally, $\textsc{ASR}(\mathcal{D}_{c},\mathcal{M})=1-\mathbb{E}_{m\in\mathcal{M}}\texttt{acc}(\texttt{LM}_{m},\mathcal{D}_{c})$.
Optimization Objective. Our goal is to find a description of harmful prompts $c$ that maximizes attack success rate, subject to a constraint that a dataset with this description $c$ exactly contains harmful prompts. We aim to solve the following constrained optimization problem:
maximize $\mathcal{J}(c;\mathcal{M})=\textsc{ASR}(\mathcal{D}_{c};\mathcal{M})$ (2)
subject to $\textsc{harm}(c)=1.$
4 Solving the Optimization Problem
We now propose an LM-based method to approximately optimize the objectives from § 3. One natural, naive design is to perform a random search, where we prompt $\texttt{LM}_{\text{evaluator}}$ to generate a diverse set of dataset descriptions $c$, prompt $\texttt{LM}_{\text{evaluator}}$ to generate a dataset of (question, answer) pairs for each description $c$, and then select the best dataset according to the objective function $\mathcal{J}(c;\mathcal{M})$.
However, this design suffers from two issues: (1) Example correctness: Since we use $\texttt{LM}_{\text{evaluator}}$ to construct the dataset, the generated answers might be incorrect due to model hallucination. (2) Example difficulty: The difficulty of the generated questions is upper-bounded by the capabilities of $\texttt{LM}_{\text{evaluator}}$ and hence cannot be used to evaluate models stronger than $\texttt{LM}_{\text{evaluator}}$. (3) Topic difficulty: empirically, in preliminary studies, we observe that $\texttt{LM}_{\text{evaluator}}$ tends to propose well-known topics, leading to insufficiently difficult dataset descriptions.
We now propose two techniques to address these issues: We first augment $\texttt{LM}_{\text{evaluator}}$ with privileged information to improve the correctness and difficulty of the generated datasets (§ 4.1). Next, we propose adaptive search, which uses the trajectory of past generated benchmarks to improve topic difficulty (§ 4.2). We present the full pseudocode of AutoBencher in Algorithm 1.